A piece of aluminum has a density of 2700 kg/m3 and a mass of 0.775 kg. The aluminum is completely submerged in a container of oil with a density of 650 kg/m3 1 A) B) What is the magnitude of the buoyant force acting on the aluminum? What would the apparent weight of the aluminum be if it were instead completely submerged in water? opparent wecighf of
Answer #1

A) Buoyant force acting on aluminium = weight of the oil displaced

= volume of aluminium x density of oil

= massof aluminium densityofaluminium densityo foil *g

\frac{0.775}{2700}*650*9.8 = 1.83 N

B) Apparent weight = weight - buoyant force

(mass of alluminium -\frac{mass of aluminium}{density of aluminium}density of water)*g

(0.775 -\frac{0.775}{2700}*1000)*g

= 4.78 N